# MAGPIE: Multi-Task Analysis of Media-Bias Generalization with Pre-Trained Identification of Expressions

### v2

@@ Abstract
Media bias detection poses a complex, multifaceted problem traditionally tackled using single-task models and small in-domain datasets, consequently lacking generalizability. To address this, we introduce MAGPIE, the first large-scale multi-task pre-training approach explicitly tailored for media bias detection. To enable pre-training at scale, we present Large Bias Mixture (LBM), a compilation of 59 bias-related tasks. MAGPIE outperforms previous approaches in media bias detection on the Bias Annotation By Experts (BABE) dataset, with a relative improvement of 3.3% F1-score. MAGPIE also performs better than previous models on 5 out of 8 tasks in the Media Bias Identification Benchmark (MBIB). Using a RoBERTa encoder, MAGPIE needs only 15% of finetuning steps compared to single-task approaches. Our evaluation shows, for instance, that tasks like sentiment and emotionality boost all learning, all tasks enhance fake news detection, and scaling tasks leads to the best results. MAGPIE confirms that MTL is a promising approach for addressing media bias detection, enhancing the accuracy and efficiency of existing models. Furthermore, LBM is the first available resource collection focused on media bias MTL.
@@
[…]
${}^{1}$zech Technical University, Prague, Czech Republic
${}^{2}$niversity of Göttingen, Göttingen, Germany, {last}@uni-goettingen.de
${}^{3}$ational Institute of Informatics, Tokyo, Japan, {last}@nii.ac.jp
${}^{4}$, TU Munich, Germany, ${}^{5}$niversity of Konstanz, Germany, {last}@uni-konstanz.de
${}^{6}${t.horych, m.wessel, t.spinde}@media-bias-research.org
@@ 1.   Introduction
Media bias is the skewed portrayal of information favoring certain group interests Recasens et al. (2013), which manifests in multiple facets, including political, gender, racial, and linguistic biases. Such subtypes of bias, which can intersect and coexist in complex combinations, make the classification of media bias a challenging task Raza et al. (2022). Existing research on media bias detection primarily involves training classifiers on small in-domain datasets Krieger et al. (2022), which exhibit limited generalizability across diverse domains Wessel et al. (2023).
@@
  • We present MAGPIE (Multi-Task Media-Bias Analysis Generalization for Pre-Trained Identification of Expressions) the first large-scale multi-task pre-training approach for media bias detection. By including diverse biases like persuasive and subjective, a classifier trained on top of MAGPIE correctly classifies sentences that state-of-the-art single-task models misidentify (we show an example in Figure 1).
@@
  • We demonstrate that MAGPIE outperforms the previous state-of-the-art model by 3.3% on the Media Bias Annotation by Experts (BABE) dataset Spinde et al. (2021c) and achieves state-of-the-art results on the Media Bias Identification Benchmark (MBIB) collection Wessel et al. (2023).
  • We make all resources, including datasets, training framework, documentation, and models, publicly available on GitHub.
@@
These contributions highlight the potential of MTL in improving media bias detection. Our findings show, e.g., that tasks like sentiment and emotionality enhance overall learning, all tasks boost fake news detection, and scaling tasks leads to optimal results. Another key insight of our research is the value of MTL in contexts where the primary dataset is small [^1] For example, the Media Bias Annotation by Experts (BABE) dataset Spinde et al. (2021c).. By learning generalized bias knowledge from a range of tasks, we can improve the accuracy and efficiency of existing models, even in the face of limited data. Overall, our research offers a novel multi-task learning approach with a variety of media bias tasks, contributing to the understanding of media bias detection.
@@ 2.1.   Media Bias
Media bias tasks and datasets mainly cover individual, self-contained tasks such as binary classifications Recasens et al. (2013); Spinde et al. (2021b), which, so far, are not explored in relation to each other Spinde et al. (2023). Wessel et al. (2023) systematically form the media bias detection benchmark MBIB by reviewing over 100 media bias datasets and consolidating 22 of them into eight distinct tasks like linguistic, racial, and political bias. Their study highlights that methods only focused on one of these tasks exhibit limitations in their detection capabilities. MAGPIE encompasses all the tasks identified in the MBIB but also significantly expands its scope by incorporating an additional 51 media bias-related tasks due to a variety of limitations in MBIB (see Section 3.1).
@@ 2.2.   Multi-Task Learning
As described above, media bias can be seen as a composite problem composed of various interrelated bias types Wessel et al. (2023). In the realm of Natural Language Understanding (NLU), MTL has proven to be highly effective when incorporating related tasks Aribandi et al. (2021b). For instance, benchmarks such as GLUE and SuperGLUE successfully decompose the NLU problem into a suite of proxy tasks, including paraphrase detection and semantic evaluation, thereby substantially improving performance across a range of NLU tasks Wang et al. (2018, 2019). Motivated by this success in NLU, we propose to jointly learn from different bias types within the media bias domain. With this approach, we aim to treat media bias not as a singular entity but as many interconnected issues.
@@ 3.   Methodolodgy
We implement MAGPIE using pre-finetuning as introduced in Aghajanyan et al. (2021) (See also Section 2). As such, MAGPIE is an encoder-only MTL transformer model pre-finetuned on 59 media bias-related auxiliary tasks provided by Large Bias Mixture (LBM), an extensive task collection of bias-related datasets [^2] LBM is also published within this paper, see Section 3.1.. We incorporate a novel approach of a Head-Specific Early Stopping and Resurrection to effectively handle tasks of varying sizes (Section 3.3.1).
@@
As outlined in Figure 2, our first step involves constructing the LBM. Following this, we define the model and multi-task learning (MTL) framework employed to train MAGPIE, which includes optimization strategies, task sampling, and auxiliary task selection. Lastly, we evaluate MAGPIE on two primary resources: the Media Bias Annotation by Experts (BABE) dataset [^3] BABE provides high-quality labels that capture a broad range of linguistic biases, thus allowing us to evaluate our model’s generalizability within a single dataset context. Spinde et al. (2021c), and the Media Bias Identification Benchmark (MBIB) collection [^4] MBIB provides a variety of bias tasks for testing and is currently the most extensive media bias benchmark available. Wessel et al. (2023).
@@ 3.1.   The LBM (Large Bias Mixture) task collection
Currently, MBIB is the only collection of media bias tasks. However, it does not include tasks that constitute a form of media bias [^5] As per the definition of directly and indirectly related media bias tasks given by Spinde et al. (2023). more indirectly, such as sentiment analysis or emotion classification. MAGPIE aims to integrate tasks both directly linked to media bias and those peripherally related, such as sentiment analysis, to provide broader coverage of linguistic features in the media bias context. Therefore, we introduce Large Bias Mixture (LBM), a more general collection of relevant media bias and media bias-related tasks, more suitable for our MTL approach. We show our task selection process in Figure 3.
First, we manually assess a list of 115 media bias-related datasets categorized into task families composed by Wessel et al. (2023). A task family is a conceptual grouping of tasks that share similar objectives, such as those related to gender bias, encompassing pronoun coreference resolution, gender classification, and sexism detection.
We use this notion of task families to analyze general knowledge transfer between media bias tasks, such as proposed by Aribandi et al. (2021b) for general language tasks.
We filter the collection of the 115 datasets based on the following criteria [^6] We base these on the criteria set by Wessel et al. (2023). However, we acknowledge that determining the dataset quality remains a manual and subjective choice.:
@@
  • Text granularity: We only use datasets labeled on a sentence level (not on, for instance, article or word level)
  • Quality of annotations: We exclude datasets with little to no documentation or unreliable annotation sources.
@@
Of the 115 datasets collected, we discard 11 datasets that are not publicly available. We discard 52 with article-level annotations and 5 with annotations on other levels [^7] One dataset has annotations on only words, two on users, and three on outlets.. We remove 5 datasets due to low quality or unreliable annotations and discard 4 duplicates. Applying these criteria leaves 38 datasets. Including 8 handpicked datasets not originally listed gives us 46 datasets. These are categorized into task families ensuring no overlap and more than two datasets per family. Finally, ten datasets with multi-level annotations, e.g., token and sentence level, are split into tasks, yielding final number of 59 tasks.
@@
The final Large Bias Mixture (LBM) includes 59 tasks, categorized into 9 distinct task families, encompassing 1,210,084 labeled sentences. We make the LBM publicly accessible, to facilitate research in media bias detection and other computational-social-science tasks. References and short descriptions of all datasets and corresponding tasks can be found in Table 4.
@@ 3.3.1.   Data sampling and early stopping
To prevent large tasks from dominating the optimization, we ensure uniform data distribution by sampling one fixed-size sub-batch from each task per training step, a regular approach in MTL Aribandi et al. (2021b); Spinde et al. (2022). We also employ early stopping as a regularization for each task individually to prevent over-fitting of tasks that converge faster. However, these methods often fall short when confronted with tasks of varied complexity and differing convergence speeds, which both is the case for tasks in LBM. When task A stops early while task B takes longer to converge, the latent representation of the shared encoder shifts toward task B. We aim to mitigate this issue by employing a training strategy that tackles the latent representation shift using two complementary approaches:
@@ 3.3.2.   Auxiliary task selection
Multi-Task Learning often involves selecting well-used datasets, leading to potential selection biases. Furthermore, manually handpicking datasets becomes challenging due to varying and potentially ambiguous bias annotation schemes. To automate the process of task selection, we utilize the GradTS algorithm Ma et al. (2021). We choose GradTS due to its demonstrated efficiency and its simplicity of implementation, which enhances its usability.
Gradient-based Task selection. In line with GradTS, we construct auxiliary tasks as follows. We individually train all tasks, accumulate absolute gradients, then extract and layer-wise normalize these in the attention heads, forming a 12x12 importance matrix [^8] RoBERTa has 12 attention heads on each of the 12 layers. for each task. Tasks are sorted by correlation between each task’s matrix and BABE task’s matrix. We pre-finetune $m-1$ models on the first $k$ tasks from the sorted list, where $k$ varies from 1 to $m-1$ and $m$ is the size of LBM. The BABE task is then evaluated on these pre-finetuned models, with the optimal $k$ determined by evaluation loss. For further details on the GradTS algorithm, please see Ma et al. (2021).
3.3.3.   Hyperparameters and training stability
Finetuning transformer models on small datasets often leads to inconsistent outcomes, such as overfitting, training instabilities, or significant performance variabilities Dodge et al. (2020). To ensure reproducibility and consistency, we use a fixed random seed for each run and perform individual task hyperparameter grid searches. To negate the influence of random weight initializations, we evaluate and compare models using 30 random seeds, reporting average performance. We use $321$ as a random seed for all experiments. For the final evaluation, where we employ 30 random seeds, we use values from $0..29$. For optimizing the model, we use a per-task batch size of 32, an AdamW optimizer, and a polynomial learning scheduler. We run all experiments on 1 NVIDIA TESLA A100 GPU with 40 GB of memory.
@@ 4.1.   Auxiliary tasks selection
Table 1: Attention importance correlation wrt media bias task.
@@ 4.2.   Evaluation
Figure 4: Final F1 score on a test set averaged over 30 random seeds. All three MTL approaches outperform baseline STL finetuning. Pre-finetuning on all LBM tasks results in significantly improved performance.
Multi-Task Learning Performance. Section 4.2 summarizes our performance results on the BABE dataset. We observe that all of our MTL pre-finetuning schemas lead to performance improvements. In particular, pre-finetuning on all tasks from LBM yields a SOTA performance on the BABE dataset, achieving an 84.1% F1 score and a relative improvement of 3.3% compared to the previous baseline by Spinde et al. (2021c). While both MTL baselines - Muppet Aghajanyan et al. (2021) and UnifiedM2 Lee et al. (2021b) outperform single-task baseline, they underperform all of our MTL models.
@@
On MBIB, our performance surpasses that of the five models evaluated by Wessel et al. (2023) in 5 out of 8 tasks. Additionally, it improves upon the current baseline with an average F1-Score of 0.7692 across all tasks. However, this improvement is only marginal. More importantly, MAGPIE outperforms the single-task RoBERTa baseline on MBIB, showing that an MTL approach is beneficial in media bias detection.

### v3

@@ Abstract
Media bias detection poses a complex, multifaceted problem traditionally tackled using single-task models and small in-domain datasets, consequently lacking generalizability. To address this, we introduce MAGPIE, a large-scale multi-task pre-training approach explicitly tailored for media bias detection. To enable large-scale pre-training, we construct Large Bias Mixture (LBM), a compilation of 59 bias-related tasks. MAGPIE outperforms previous approaches in media bias detection on the Bias Annotation By Experts (BABE) dataset, with a relative improvement of 3.3% F1-score. Furthermore, using a RoBERTa encoder, we show that MAGPIE needs only 15% of fine-tuning steps compared to single-task approaches. We provide insight into task learning interference and show that sentiment analysis and emotion detection help learning of all other tasks, and scaling the number of tasks leads to the best results. MAGPIE confirms that MTL is a promising approach for addressing media bias detection, enhancing the accuracy and efficiency of existing models. Furthermore, LBM is the first available resource collection focused on media bias MTL.
@@
Tomáš Horych1,3,6, Martin Wessel4,6, Jan Philip Wahle2, Terry Ruas2,
Jerome Waßmuth5, André Greiner-Petter2,3, Akiko Aizawa3,
Bela Gipp2, Timo Spinde2,6
1Czech Technical University, Prague, Czech Republic
2University of Göttingen, Göttingen, Germany, {last}@uni-goettingen.de
3National Institute of Informatics, Tokyo, Japan, {last}@nii.ac.jp
4CDTM, TU Munich, Germany, 5University of Konstanz, Germany, {last}@uni-konstanz.de
6{t.horych, m.wessel, t.spinde}@media-bias-research.org
@@ 1.   Introduction
Media bias is a skewed portrayal of information favoring certain group interests Recasens et al. (2013), which manifests in multiple facets, including political, gender, racial, and linguistic biases. Such subtypes of bias, which can intersect and coexist in complex combinations, make the classification of media bias a challenging task Raza et al. (2022). Existing research on media bias detection primarily involves training classifiers on small in-domain datasets Krieger et al. (2022), which exhibit limited generalizability across diverse domains Wessel et al. (2023).
@@
  1. 1.
    We present MAGPIE - the first large-scale multi-task pre-training approach for media bias detection. By pre-training on diverse bias types such as persuasive and subjective, a classifier based on MAGPIE correctly classifies sentences that state-of-the-art single-task models misidentify (we show an example in Figure 1).
  2. 2.
@@
  4. 4.
    We demonstrate that MAGPIE outperforms the previous state-of-the-art model by 3.3% on the Media Bias Annotation by Experts (BABE) dataset Spinde et al. (2021c) and achieves competetive results on the Media Bias Identification Benchmark (MBIB) collection Wessel et al. (2023).
  5. 5.
    We make all resources, including datasets, training framework, documentation, and models, publicly available on GitHub:
@@
These contributions highlight the potential of MTL in improving media bias detection. Our findings show, e.g., that tasks like sentiment and emotionality enhance overall learning, all tasks boost fake news detection, and scaling tasks leads to optimal results. Another key insight of our research is the value of MTL in contexts where the primary dataset is small [^1] For example, the Media Bias Annotation by Experts (BABE) dataset Spinde et al. (2021c).. By learning generalized bias knowledge from a range of tasks, we can improve the accuracy and efficiency of existing models, even in the face of limited data. Overall, our research offers a multi-task learning approach to media bias detection with first large-scale resources in the domain.
@@ 2.1.   Media Bias
Media bias tasks and datasets mainly cover individual, self-contained tasks such as binary classifications Recasens et al. (2013); Spinde et al. (2021b), which, so far, are not explored in relation to each other Spinde et al. (2023). Wessel et al. (2023) systematically form the media bias detection benchmark MBIB by reviewing over 100 media bias datasets and consolidating 22 of them into eight distinct tasks like linguistic, racial, and political bias. Their study highlights that methods only focused on one of these tasks exhibit limitations in their detection capabilities. MAGPIE encompasses all the tasks identified in the MBIB but also significantly expands its scope by incorporating an additional 51 media bias-related tasks to mitigate a variety of limitations in MBIB (see Section 3.1).
@@ 2.2.   Multi-Task Learning
As described above, media bias can be seen as a composite problem composed of various interrelated bias types Wessel et al. (2023). In the realm of Natural Language Understanding (NLU), MTL has proven to be highly effective when incorporating related tasks Aribandi et al. (2021b). For instance, benchmarks such as GLUE and SuperGLUE successfully decompose the NLU problem into a suite of proxy tasks, including paraphrase detection Wahle et al. (2023, 2022), and semantic evaluation Ruas et al. (2019, 2020), thereby substantially improving performance across a range of NLU tasks Wang et al. (2018, 2019). Motivated by this success in NLU, we propose to jointly learn from different bias types within the media bias domain. With this approach, we aim to treat media bias not as a singular entity but as many interconnected issues.
@@ 3.   Methodolodgy
We implement MAGPIE using pre-finetuning as introduced in Aghajanyan et al. (2021) (See also Section 2). As such, MAGPIE is an encoder-only MTL transformer model pre-finetuned on 59 media bias-related auxiliary tasks provided by Large Bias Mixture (LBM), a large-scale task collection of bias-related datasets. We incorporate a novel approach of a Head-Specific Early Stopping and Resurrection to effectively handle tasks of varying sizes (Section 3.3.1).
@@
As outlined in Figure 2, our first step involves constructing the LBM. Following this, we define the model and multi-task learning (MTL) framework employed to train MAGPIE, which includes optimization strategies, task sampling, and auxiliary task selection. Lastly, we evaluate MAGPIE on two primary resources: the Media Bias Annotation by Experts (BABE) dataset [^2] BABE provides high-quality labels that capture a broad range of linguistic biases, thus allowing us to evaluate our model’s generalizability within a single dataset context. Spinde et al. (2021c), and the Media Bias Identification Benchmark (MBIB) collection.
@@ 3.1.   The LBM (Large Bias Mixture) task collection
Currently, MBIB is the only collection of media bias tasks. However, it does not include tasks that constitute a form of media bias more indirectly, such as sentiment analysis or emotion classification. MAGPIE aims to integrate tasks both directly linked to media bias and those peripherally related, such as sentiment analysis, to provide broader coverage of linguistic features in the media bias context. Therefore, we introduce Large Bias Mixture (LBM), a more general collection of relevant media bias and media bias-related tasks, more suitable for our MTL approach. We show our task selection process in Figure 3.
First, we manually assess a list of 115 media bias-related datasets in English language, categorized into task families by Wessel et al. (2023). A task family is a conceptual grouping of tasks that share similar objectives, such as those related to gender bias, encompassing pronoun coreference resolution, gender classification, and sexism detection.
We use this notion of task families to analyze general knowledge transfer between media bias tasks in Section 4.3, such as Aribandi et al. (2021b) proposed for general NLP tasks.
We filter the collection of the 115 datasets based on the following criteria [^3] We acknowledge that determining the dataset quality remains a manual and subjective choice.:
@@
  • Text granularity: We only use datasets labeled on a sentence level or its fragments (tokens)(not on, for instance, article level)
  • Quality of annotations: We exclude datasets with no documentation, low annotation agreement or employment of machine annotation.
@@
Of the 115 datasets collected, we discard 11 datasets that are not publicly available. We discard 52 with article-level annotations and 5 with annotations on other levels [^4] One discarded dataset provides only a list of biased words, other two annotations on users, and three on outlets.. We remove 5 datasets due to or unreliable source of annotations and discard 4 duplicates. Applying these criteria leaves 38 datasets. Including 8 handpicked datasets not originally listed gives us 46 datasets. These are categorized into task families ensuring no overlap and more than two datasets per family. Finally, ten datasets with multi-level annotations, e.g., token and sentence level, are split into tasks, yielding final number of 59 tasks.
@@
The final Large Bias Mixture (LBM) includes 59 tasks, categorized into 9 distinct task families, encompassing 1,210,084 labeled sentences. We make the LBM publicly accessible, to facilitate research in media bias detection and other computational-social-science tasks. References and short descriptions of all datasets and corresponding tasks and task families can be found in Table 4.
@@ 3.3.1.   Data sampling and early stopping
To prevent large tasks from dominating the optimization, we ensure uniform data distribution by sampling one fixed-size sub-batch from each task per training step, a regular approach in MTL Aribandi et al. (2021b); Spinde et al. (2022). We also employ early stopping as a regularization for each task individually to prevent over-fitting of tasks that converge faster. However, these methods often fall short when confronted with tasks of varied complexity and differing convergence speeds, which both is the case for tasks in LBM. When task A stops early while task B takes longer to converge, the latent representation of the shared encoder shifts toward task B.
We aim to mitigate this issue by employing a training strategy that tackles the latent representation shift using two complementary approaches:
@@ 3.3.2.   Auxiliary task selection
Multi-Task Learning often involves selecting well-used datasets, leading to potential selection biases. Furthermore, manually handpicking datasets becomes challenging due to varying and potentially ambiguous bias annotation schemes.
To automate the process of task selection, we utilize the GradTS algorithm Ma et al. (2021). We choose GradTS due to its demonstrated efficiency and its simplicity of implementation, which enhances its usability.
Gradient-based Task selection. In line with GradTS, we construct auxiliary tasks as follows. We individually train all tasks, accumulate absolute gradients, then extract and layer-wise normalize these in the attention heads, forming a 12x12 importance matrix [^5] RoBERTa has 12 attention heads on each of the 12 layers. for each task. Tasks are sorted by correlation between each task’s matrix and BABE task’s matrix.
We pre-finetune $m-1$ models on the first $k$ tasks from the sorted list, where $k$ varies from 1 to $m-1$ and $m$ is the size of LBM. The BABE task is then evaluated on these pre-finetuned models, with the optimal $k$ determined by evaluation loss. For further details on the GradTS algorithm, please see Ma et al. (2021).
3.3.3.   Experimental setup
We split the BABE dataset into train, dev and test split with 70%,15%,15% portions respectively. For both hyperparameter tuning and auxiliary task selection, we evaluate on dev split and only use test split for the final evaluation (Section 4.2).
As fine-tuning transformers on small datasets often leads to inconsistent results, such as high performance variance Dodge et al. (2020), we use a fixed random seed $321$ for all runs in auxilliary task selection.
For the final evaluation on the test set (Table 2) we evaluate the models using 30 random seeds and report the average performance, to minimize the influence of random weight initializations. We use values $0..29$.
For optimizing the models, we use a per-task batch size of 32, an AdamW optimizer, and a polynomial learning scheduler. We run all experiments on 1 NVIDIA TESLA A100 GPU with 40 GB of memory.
@@ 4.1.   Auxiliary tasks selection
Table 1: Attention importance correlation w.r.t the media bias task.
@@ 4.2.   Evaluation
Figure 4: Final F1 score on a BABE test set averaged over 30 random seeds. All three MTL approaches outperform baseline STL finetuning. Pre-finetuning on all LBM tasks results in significantly improved performance.
Multi-Task Learning Performance. Table 2 summarizes our performance results on the BABE dataset. We observe that all of our MTL pre-finetuning schemas lead to performance improvements. In particular, pre-finetuning on all tasks from LBM yields a SOTA performance on the BABE dataset, achieving an 84.1% F1 score and a relative improvement of 3.3% compared to the previous baseline by Spinde et al. (2021c). While both MTL baselines - Muppet Aghajanyan et al. (2021) and UnifiedM2 Lee et al. (2021b) outperform single-task baseline, they underperform all of our MTL models.
On MBIB benchmark, MAGPIE ranks first on 5 out of 8 tasks. However, the improvements are only marginal. The results can be found in Appendix in Table 5.
@@ Appendix A Appendix
MBIB Task MAGPIE RoBERTa ConvBERT
Linguistic Bias 0.7139 0.7076 0.7126
Cognitive Bias 0.7086 0.7037 0.7044
Text-Level Context Bias 0.7638 0.7646 0.7697
Hatespeech 0.8747 0.8759 0.8805
Gender Bias 0.8344 0.8322 0.8257
Racial Bias 0.8809 0.8761 0.8772
Fake News 0.6709 0.6711 0.6787
Political Bias 0.7059 0.7029 0.7041
Table 5: Performance of MAGPIE and two baselines RoBERTa, and ConvBERT (state-of-the-art model in Wessel et al. (2023)) on the Media Bias Identification Benchmark (MBIB) tasks.